# Frank–Wolfe algorithm for DC optimization problem

### v2

@@ 5 Classical Frank–Wolfe algorithm with an adaptive stepsize
Let $\epsilon>0$. Then, Algorithm 1 generates a point $x^{k}$ such that $|{\omega}_{k}|\leq\epsilon$, performing at most,
$\left(\max\left\{{\Gamma}^{2},{\alpha\Gamma}\right\}\frac{1}{\epsilon^{2}}\right)\left(2+\log_{2}\left({\beta}\frac{1}{\epsilon^{2}}\right)\right)={\cal O}(|\log_{2}(\epsilon)|\epsilon^{-2})$
evaluations of functions $f$, and $\max\left\{{\Gamma}^{2},{\alpha\Gamma}\right\}\frac{1}{\epsilon^{2}}+1={\cal O}(\epsilon^{-2})$ evaluations of gradients of $f$ and subgradients of $h$.
5.1 Iteration-complexity analysis: The piecewise weak-star-convex case
In this section, we establish iteration–complexity bounds for the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 when the objective function is a DC function f=ghf=g-h that is piecewise weak–star–convex, where the first component $g$ has a Lipschitz continuous gradient on ${\cal C}\subset{\mathbb{R}}^{n}$ and the second component $h$ is convex but not necessarily differentiable. As will become clear, the piecewise-star-convexity property plays a central role in our analysis. By combining this property with the descent estimates from Proposition 5.1, we obtain the following global complexity result for the sequence $(x^{k})_{k\in\mathbb{N}}$ generated by Algorithm 1.
Theorem 5.5.
Assume that $f:\mathbb{R}^{n}\to\mathbb{R}$ is a piecewise-star-convex on $\cal C$}, and let ${\cal P}=\{\,{\cal C}_{i}:i\in I\}$ be a partition of $\cal C$} into nonempty convex subsets such that $\bigcup_{i\in I}{\cal C}_{i}={\cal C}$. For each $i\in I$, let $X_{i}^{*}$ denote the set of global minimizers of $f$ on ${\cal C}_{i}$, and let
@@ 5.1 Iteration-complexity analysis: The piecewise weak-star-convex case
be the corresponding minimum value. Let $(x^{k})_{k\in\mathbb{N}}$ be the sequence generated by Algorithm 1. Then there exists $i_{0}\in I$ such that:
$f(x^{k})-f_{i_{0}}^{*}\leq\displaystyle\frac{4(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k}$, for all $k=1,2,\ldots.$
$\min_{\ell\in\left\{\lfloor\frac{k}{2}\rfloor+2,...,k\right\}}|{\omega}_{\ell}|\leq\frac{16(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k-2}$, for all $k=3,4,\ldots,$ where $\lfloor k/2\rfloor=\max_{n\in\mathbb{N}}\left\{n\leq k/2\right\}.$
@@
By (28) in Proposition 5.1, we have
$f\!\big{(}x^{k}+\lambda_{k}(p^{k}-x^{k})\big{)}\leq f(x^{k})-|\omega(x^{k})|\,\lambda_{k}\;+\;\frac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\lambda_{k}^{2}.$ (36)
On the other hand, by (21),
$\lambda_{k}\;=\;\operatorname*{arg\,min}_{\lambda\in(0,1]}\Big{\{}-|\omega(x^{k})|\,\lambda+\tfrac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\lambda^{2}\Big{\}}.$
Hence, for any choice $\beta_{k}\in(0,1]$, the right-hand side of (36) at $\lambda=\lambda_{k}$ is bounded by its value at $\lambda=\beta_{k}$. Taking $\beta_{k}:=\tfrac{2}{k+2}$, we get
$f\!\big{(}x^{k}+\lambda_{k}(p^{k}-x^{k})\big{)}\leq f(x^{k})-|\omega_{k}|\,\beta_{k}\;+\;\frac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\beta_{k}^{2}.$
Using $\|p^{k}-x^{k}\|\leq\operatorname{diam}({\cal C})$, together with (23) and Lemma 5.2 (which yields $\tfrac{2^{\,j_{k}}L_{k}}{2}\leq L+L_{0}$), we obtain
@@
Since $\cal C$} is compact, the sequence $(x^{k})_{k\in\mathbb{N}}$ generated by Algorithm 1 is bounded. Let $\bar{x}$ be a limit point, and let $(x^{k_{\ell}})_{\ell\in\mathbb{N}}$ be a subsequence such that $\lim_{\ell\to+\infty}x^{k_{\ell}}={\bar{x}}$. Because ${\cal P}=\{{\cal C}_{i}:i\in I\}$ is a partition of $\cal C$}, there exists $i_{0}\in I$ with $\bar{x}\in{\cal C}_{i_{0}}$. In addition, Theorem 5.3 implies that $\lim_{k\to+\infty}{\omega}_{k}=0$ and $\bar{x}$} is a stationary point to problem (14). Hence, by (3.12), $\bar{x}$ is a minimizer of $f$ on ${\cal C}_{i_{0}}$. Let $f_{i_{0}}^{*}$ denote the minimum value of $f$ on ${\cal C}_{i_{0}}$. Thus, $\lim_{\ell\to+\infty}f(x^{k_{\ell}})=f({\bar{x}})=f_{i_{0}}^{*}$. Moreover, (29) in Proposition 5.1 implies that $\big{(}f(x^{k})\big{)}_{k\in\mathbb{N}}$ is monotone nonincreasing. Therefore,
$\lim_{k\to\infty}f(x^{k})\;=\;f_{i_{0}}^{*},\qquad\text{and}\qquad f_{i_{0}}^{*}\leq f(x^{k})\quad\forall\,k\in\mathbb{N}.$ (38)
@@
Since $f$ is piecewise weak–star–convex on $\cal C$}, Proposition 3.11 and Theorem 2.2 ensure that, for any $x^{\star}\in X_{i_{0}}^{*}$,
$f_{i_{0}}^{*}-f(x^{k})\geq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(x^{*}_{k}-x^{k}).$
where $u^{k}\in\partial h(x^{k})$. Using (20) and (26), the definition of $\omega_{k}$ gives $0\geq f_{i_{0}}^{*}-f(x^{k})\geq\omega_{k},$ hence
$0\leq f(x^{k})-f_{i_{0}}^{*}\leq|\omega_{k}|\qquad\forall\,k\in\mathbb{N}.$ (39)
Now apply Lemma 2.8 to the sequences
$a_{k}=f(x^{k})-f_{i_{0}}^{*}\leq b_{k}=|\omega_{k}|,\qquad\beta_{k}=\frac{2}{k+2},\qquad A=(L+L_{0})\,\operatorname{diam}({\cal C})^{2}$

### v3

@@ 5.2 Iteration-complexity analysis
Let ${\cal C}\subset\mathbb{R}^{n}$ be compact and convex, and let ${\cal P}=\{{\cal C}_{i}:i=1,\dots,m\}$ be a finite partition of $\cal C$} into nonempty convex sets. Assume that $f:\mathbb{R}^{n}\to\mathbb{R}$ is piecewise-star-convex on $\cal C$}. For each $i\in\{1,\dots,m\}$, let $X_{i}^{*}$ denote the set of global minimizers of $f$ on ${\cal C}_{i}$, and let
@@
be the corresponding minimum value. Let $\{x^{k}\}_{k\in\mathbb{N}}$ be the sequence generated by Algorithm 1. Then there exist $i_{*}\in\{1,\dots,m\}$ and $N\in{\mathbb{N}}$ such that:
  • (i)
    $f(x^{k})-f_{i_{*}}^{*}\leq\displaystyle\frac{4(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k-N+1}$, for all $k\geq N$;
  • (ii)
    $\min_{\ell\in\left\{\lfloor{k-N+1}/{2}\rfloor+N+1,...,k\right\}}|{\omega}_{\ell}|\leq\frac{16(L+L_{0})\operatorname{diam}({\cal C})^{2}}{k-N-1}$, for all $k\geq N+2$.
@@
Using (22) in Proposition 5.1 and recalling that $\omega_{k}\leq 0$, we obtain that
$f\!\big(x^{k}+\lambda_{k}(p^{k}-x^{k})\big)\leq f(x^{k})-|\omega_{k}|\lambda_{k}+\frac{2^{\,j_{k}}L_{k}}{2}\|p^{k}-x^{k}\|^{2}\,\lambda_{k}^{2}.$ (32)
On the other hand, by definition of the Lipschitz-based adaptive stepsize $\lambda_{k}$, we have
$\lambda_{k}=\operatorname*{arg\,min}_{\lambda\in(0,1]}\Big\{-|\omega_{k}|\,\lambda+\tfrac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\lambda^{2}\Big\}.$
Hence, for any $\beta_{k}\in(0,1]$, in particular for $\beta_{k}:=\tfrac{2}{k+2}$, we conclude from the last equality that
$|\omega_{k}|\,\lambda_{k}+\frac{2^{\,j_{k}}L_{k}}{2}\|p^{k}-x^{k}\|^{2}\,\lambda_{k}^{2}\leq-|\omega_{k}|\,\beta_{k}+\tfrac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\beta_{k}^{2}.$
Thus, combining this previous inequality with the inequality in (32), we conclude that
$f(x^{k+1})\leq f(x^{k})-|\omega_{k}|\,\beta_{k}+\frac{2^{\,j_{k}}L_{k}}{2}\,\|p^{k}-x^{k}\|^{2}\,\beta_{k}^{2}.$
Since $\|p^{k}-x^{k}\|\leq\operatorname{diam}({\cal C})$ and $\frac{2^{\,j_{k}}L_{k}}{2}=L_{k+1}\leq L+L_{0}$, by (19) and Lemma 5.2, we obtain
@@
Set $A:=(L+L_{0})\,\operatorname{diam}({\cal C})^{2}$. Since $\cal C$} is compact, $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded and has accumulation points. By Theorem 5.3(i), $\{f(x^{k})\}_{k\in\mathbb{N}}$ is monotone nonincreasing and thus convergent. Let $\bar{x}$ be a limit point, and let $(x^{k_{\ell}})_{\ell\in\mathbb{N}}$ be a subsequence such that $\lim_{\ell\to+\infty}x^{k_{\ell}}={\bar{x}}$. By Theorem 5.3(iii), there exists an index $i_{*}\in\{1,\dots,m\}$ such that $\bar{x}\in{\cal C}_{i_{*}}$ and $\bar{x}\in X_{i_{*}}^{*}$. Hence, we have
$\lim_{k\to\infty}f(x^{k})=\lim_{\ell\to\infty}f(x^{k_{\ell}})=f(\bar{x})=f_{i_{*}}^{*},\qquad\text{and}\qquad f_{i_{*}}^{*}\leq f(x^{k})\ \ \forall\,k\in\mathbb{N}.$
Now applying Lemma 5.4 to the monotone sequence $\{f(x^{k})\}$ and the finite partition $\cal P$}.we conclude that there exists $N\in\mathbb{N}$ such that, for every $k\geq N$ and every index $i_{k}$ with $x^{k}\in{\cal C}_{i_{k}}$,
$f_{i_{k}}^{*}\leq f_{i_{*}}^{*}.$ (34)
@@
Fix $k$ and choose $i_{k}$ with $x^{k}\in{\cal C}_{i_{k}}$. Select $x_{k}^{*}\in X_{i_{k}}^{*}$ such that $f(x_{k}^{*})=f_{i_{k}}^{*}$. Since $f$ is piecewise-star-convex on $\cal C$}, Proposition 4.9 and Theorem 2.2 yield
$f_{i_{k}}^{*}-f(x^{k})\geq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(x^{*}_{k}-x^{k}),$ (35)
for some $u^{k}\in\partial h(x^{k})$. Combining (34) with (35), we obtain, for all $k\geq N$, that
$f_{i_{*}}^{*}-f(x^{k})\geq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(x^{*}_{k}-x^{k}).$ (36)
Moreover, since $p^{k}$ solves the LO oracle (16), we conclude that for every $x\in{\cal C}$, there holds
$(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(x-x^{k})\ \geq\ (\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{k}-x^{k})=\omega_{k}.$
Applying this with $x=x_{k}^{*}$ and using (36), we obtain $0\geq f_{i_{*}}^{*}-f(x^{k})\geq\omega_{k}$ for all $k\geq N$. Hence,
$0\leq f(x^{k})-f_{i_{*}}^{*}\leq|\omega_{k}|\qquad\forall\,k\geq N.$ (37)
For each integer ${m}\geq 1$, define
$a_{m}:=f\!\big(x^{N+{m}-1}\big)-f_{i_{*}}^{*},\qquad b_{m}:=\big|\omega_{\,N+{m}-1}\big|,\qquad\beta_{m}:=\frac{2}{{m}+1}.$ (38)
Applying (33) with $A:=(L+L_{0})\,\operatorname{diam}({\cal C})^{2}$, k=N+m1k=N+{m}-1 and $\beta_{k}=\beta_{m}$, we get for all ${m}\geq 1$,
$a_{{m}+1}\leq a_{m}-b_{m}\,\beta_{m}+A\,\beta_{m}^{2}.$